# Garf: Learning Generalizable 3D Reassembly for Real-World Fractures

### v1

@@ 4.2 Can Garf Generalize to Unseen Shapes?
Results and Analysis. Table 2 presents the evaluation results on the Breaking Bad dataset, where Garf achieves significant improvements over previous SOTA methods. Compared to PF++ [51], Garf reduces rotation error by 82.87% and translation error by 79.83%, while achieving a CD below 0.001, indicating near-perfect reconstructions. Even more impressively, Garf-mini demonstrates exceptional generalization capability. Despite being trained only on the Everyday subset, it maintains consistent performance on the unseen Artifact subset, avoiding the performance degradation typically observed with unseen object shapes. This highlights Garf’s robust feature extraction and assembly mechanisms. Notably, Garf achieves 95.33% and 95.04% PA on the Breaking Bad dataset, approaching the theoretical maximum PA of 96.49% (Everyday subset) and 96.10% (Artifact subset) when excluding fragments smaller than 0.1% of the object volumeScanning tiny 3D fragments in real-world settings is inherently challenging; anthropologists often treat such fragments as missing parts.. On the challenging synthetic subset of Fractura, which contains unseen domain-specific objects, Garf further demonstrates superior generalization capability, outperforming all competing methods across all evaluation metrics. As shown in Fig. 4, Garf consistently produces more accurate reassemblies, particularly on the Breaking Bad Artifact subset and the Fractura synthetic fracture subset, confirming its strong generalization to unseen object shapes.
Table 2: Quantitative Results on Volume-Constrained Breaking Bad [42] and Fractura Datasets. The best performance metric is highlighted in bold, while the second-best is underlined.
@@
degree ×$\times 10^{-2}$$\times 10^{-3}$
@@
Global [22] 80.50 14.60 28.70 13.00
LSTM [53] 82.70 15.10 27.50 13.30
DGL [60] 80.30 13.90 31.60 11.80
Jigsaw [30] 42.19 6.85 68.89 8.22
PMTR [21] 31.57 9.95 70.60 5.56
PF++ [51] 35.61 6.05 76.17 2.78
@@ 4.3 How Do Fracture Types Affect Generalization?
As scanning time significantly increases with fragment count [18], real fracture datasets are typically limited in size and diversity, making large-scale training infeasible. To address this, we investigate how fracture types impact zero-shot generalization on Fantastic Breaks and Fractura, as well as the fine-tuning performance using domain-specific synthetic fractures on Fractura.
Real Fracture on Fantastic Breaks [18]. As shown in Table 3, Garf demonstrates superior generalization to real-world fracture surfaces, achieving a remarkable 48.65% reduction in rotation error compared to PF++ [51], indicating that our model effectively bridges the synthetic-to-real gap in fracture surface understanding for everyday objects.
Real Fracture on Fractura. We further evaluate performance across three fracture types. To isolate the impact of unseen objects, $\textsc{Garf}{}_{\text{LoRA}}$ is fine-tuned separately on synthetic fractures from bones, eggshells, and lithics in FracturaFine-tuning is not performed for ceramics, as its object categories closely resemble those in the Everyday subset of Breaking Bad.. Figure 5 compares reassembly performance across three fracture types in the Fractura real fracture subset. Garf outperforms competing methods on ceramics even with three missing fragments. $\textsc{Garf}{}_{\text{LoRA}}$ further improves performance by mitigating the effect of unseen objects. $\textsc{Garf}{}_{\text{LoRA}}$ generalize well to random breakage (limb bones, ceramics) and incomplete ossification (vertebrae). Unfortunately, although finetune significantly improves the performance on lithics, both Garf and $\textsc{Garf}{}_{\text{LoRA}}$ struggle on lithics, likely due to high ambiguity among flakes and the core, a well-known challenging spatial reasoning problem for anthropologists [20]. This unresolved challenge in Fractura offer valuable directions for future research.
Table 3: Quantitative Results on Fantastic Breaks Dataset [18]. This includes manually collected real-world objects.
@@
degree ×$\times 10^{-2}$$\times 10^{-3}$
@@ 4.5 Ablation Study
Pretraining Strategy. We evaluate our fracture-aware pretraining strategy by comparing PF++ [51] with two variants using the same diffusion model: (i) VAE-based pretraining [48], and (ii) our fracture-aware pretraining. As shown in Table 5, our strategy reduces RMSE(R) by 53.33% and RMSE(T) by 49.05% compared to the VAE-based strategy.
Table 5: Ablation Study on the Fracture-Aware Pretraining.
RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$ CD $\downarrow$
PF++ (VAE) 32.91 5.26 78.95 3.04
PF++ (Fracture-aware) 15.36 2.68 89.40 0.66
Table 6: Ablation Study on Our Designs of FM.
[…]
Designs in FM. We conduct an ablation study to evaluate the impact of design choices in our FM module. As shown in Table 6, vanilla FM, trained with spherical linear interpolation (slerp) to approximate valid rotations in the forward process [12], achieves 89.08 PA, already surpassing previous methods [51, 30]. Incorporating the $\mathrm{SE}(3)$ representation further improves performance by pre-modeling the manifold distribution and better capturing distribution shifts during assembly. Multi-anchor training strategy further enhances results, while one-step pre-assembly significantly boosts performance by providing a more reasonable initial pose distribution, leading to the best overall outcomes.
Sample Steps / One-Step Initialization. Table 7 shows the effect of varying sampling steps in our FM framework. Surprisingly, even with just 5 steps, FM achieves 93.70% PA, highlighting its effectiveness in modeling global probabilistic paths. Additionally, our one-step pre-assembly provides a more reasonable initial pose, further improving assembly quality while adding minimal computational overhead.
Anchor Fragment. Similar to PF++ [51], we use the largest fragment as the anchor fragment at inference. We compare the performance of using the largest fragment, a randomly selected fragment, and no anchor fragment. As shown in Table 8, using a random fragment as the anchor fragment has almost no negative effect on the model. Only anchor-free initialization leads to a slight performance drop.
Comparison with Diffusion Models. Table 9 compares our FM module with diffusion models. While diffusion, when paired with fracture-aware pretraining, achieves competitive performance, directly applying vanilla FM yields lower results (89.08 PA), emphasizing the importance of our subsequent design choices. A key limitation of diffusion models is their handling of $\mathrm{SO}(3)$ rotation, which cannot be naturally incorporated into the reverse process. Existing methods, such as score prediction [59], aim to maintain rotation validity but fall outside our current scope. Additionally, diffusion models rely on multi-step denoising without explicitly modeling the global probabilistic path, rendering one-step pre-assembly ineffective. Furthermore, on Fractura, diffusion models exhibit weaker generalization to unseen objects compared to Garf-mini.
Figure 7: Performance vs. Fragment Count on the Artifact Subset. Garf-mini significantly outperforms PF++, with one-step pre-assembly further boosting results for $>10$ fragments.
Number of Fragments. We analyze performance across varying fragment counts, as shown in Figure 7. Garf-mini consistently surpasses PF++ across all fragment counts. Our one-step pre-assembly further enhances performance on unseen objects containing more than 10 fragments.
@@ 5 Impact and Limitations
Limitations and Future Directions. Garf encounters challenges in handling fracture ambiguity, especially when fragments have subtle geometric differences or inherently ambiguous fracture surfaces. For instance, it struggles with lithic refitting in Fractura due to high ambiguity among flakes and the core, as well as fresco reconstruction in RePAIR, where erosion affects fracture surfaces [47]. Therefore, our future work will focus on: (i) Multimodal fracture reassembly, integrating geometric and texture information; (ii) Test-time policy optimization using expert feedback; (iii) Expanding the size and diversity of Fractura.
@@ 6 Related Work
Table 7: Ablation Study on Sample Steps.
1 12.52 3.18 86.88 2.14 38.26
One-Step + 1 9.79 2.46 91.31 1.42 45.76
5 8.25 1.92 93.70 0.53 57.32
One-Step + 5 7.15 1.66 94.43 0.46 76.23
20 7.63 1.60 94.02 0.35 185.05
One-step + 20 6.68 1.34 94.77 0.25 190.77
50 7.50 1.54 94.01 0.32 408.40
Table 8: Ablation Study on the Different Anchor Initialization.
Largest Anchor 6.10 1.22 95.33 0.22
Random Anchor 6.09 1.30 95.20 0.29
Anchor-Free 9.09 2.13 93.23 0.91
Table 9: Comparison Between Diffusion and Our FM Models.
Dataset Methods RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
Everyday Diffusion 7.45 1.47 94.30
$\mathrm{SE}(3)$ Diffusion N/A N/A N/A
Diffusion w/ One-Step 7.51 1.47 94.27
Vanilla FM 10.24 1.95 89.08
Garf-mini 6.68 1.34 94.77
Fractura Diffusion $32.38$ $7.90$ $71.73$
Garf-mini 27.88 6.79 76.25
Fracture Assembly. Fracture Assembly is a challenging spatial intelligence task. Early methods relied on explicit geometric matching with handcrafted features [31, 2, 56, 14], often struggling with ambiguous or incomplete geometries. The advent of the large-scale synthetic dataset Breaking Bad [42] has enabled learning-based approaches to acquire robust geometric representations and assembly strategies [29]. Jigsaw [30] jointly learns hierarchical features from global and local geometries for fracture matching and pose estimation, while SE(3)-equiv [54] extracts fragment features for pose estimation. DiffAssemble [41] improves performance using a diffusion model. PuzzleFusion++ [51] mimics how humans solve spatial puzzles by integrating diffusion-based pose estimation with a VAE-based fragment representation and transformer-based alignment verification. However, while PuzzleFusion++ achieves SOTA results on the everyday subset, its performance degrades significantly on the Artifact subset [51]. More critically, it remains unclear whether models trained on synthetic data can generalize to real-world fractures with more complex breakage patterns. To fill this gap, we identify major real-world fracture challenges and curate Fractura, a dataset capturing key complexities. To address these challenges, we propose Garf, a generalizable 3D reassembly framework for real-world fractures.
Flow Matching. Flow matching (FM) has emerged as a powerful alternative to diffusion models, offering advantages such as simulation-free training, closed-form target vector fields, and more efficient optimization [27, 25, 38]. Unlike diffusion models that require iterative denoising, FM directly learns a vector field that smoothly transforms a prior distribution into the target distribution along straight probability paths. Recent advances have explored FM across image generation [9], protein backbone generation [58, 16, 11], and general robot control [3], demonstrating superior efficiency compared to diffusion-based alternatives. AssembleFlow [12] attempts to leverage FM for molecular assembly, but introduces numerical errors by approximating quaternion updates through direct addition over small time intervals during inference. This limitation has been analyzed in our ablation study (Section 4.5). While diffusion models have been widely applied to fracture assembly [41, 51], FM provides a more natural formulation by learning geodesic flows in SE(3). To leverage these advantages, we propose the first FM-based fracture assembly framework, incorporating a multi-anchor training strategy and a one-step pre-assembly at inference time.
@@ 7 Conclusion
Acknowledgement We gratefully acknowledge the Physical Anthropology Unit, Universidad Complutense de Madrid for access to curated human skeletons, and Dr. Scott A. Williams (NYU Anthropology Department) for the processed data samples. This work was supported in part through NSF grants 2152565, 2238968, 2322242, and 2426993, and the NYU IT High Performance Computing resources, services, and staff expertise.
@@ A.1 Fracture Simulation
(i) Bone. For elongated structures like limbs and ribs, we used Blender’s skinning and subdivision surface techniques to create realistic cylindrical hollows, replicating bone morphology. We then applied the physics-based fracture method from Breaking Bad [42] to generate 2–20 fragments. The same approach was used for os coxae and vertebrae, forming the simulated subset of the bone category. (ii) Eggshell. Since scanned eggshells produce watertight solid ellipsoids, we removed 98% of the concentric volume to simulate thin shells. We then applied the same physics-based fracture method to generate realistic breakage patterns. (iii) Ceramics. Given that ceramic objects (e.g., bowls, pots, vases) closely resemble those in Breaking Bad’s everyday category, we focused on scanning real fragments and did not include a simulated subset. (iv) Lithics. As an initial feasibility test, two generalized core morphologies were repeatedly virtually knapped with some randomized variation following methods described for the dataset in [36] to produce core and flake combinations with varying geometries.
@@ B.1 Data Preprocessing
We preprocess the BreakingBad dataset [42] to calculate the segmentation ground truth directly from meshes to reduce the computation overhead during training as described in Sec. 3.1, and there’s no need for any hyperparameters. Unlike baseline methods (Global, LSTM, and DGL) provided by the dataset and PF++ [51], which samples $M=1000$ points from the mesh per fragment, we used the same setting as in Jigsaw [30] to sample $M=5000$ points per object, making all fragments have the same point density. With this sampling setting, we did not encounter any gradient explosion issues during training, as reported in FragmentDiff [57], which occur when sampling too many points for tiny pieces. Meanwhile, we employ the Poisson disk sampling method to ensure that the points are more uniformly distributed on the surface of the fragment. During training, standard data augmentation techniques are applied, including recentering, scaling, and random rotation.
@@ B.2 Training Recipe
We modified a smaller version of Point Transformer V3 [55] as our backbone for the segmentation pretraining, as shown in Table II, which we found to be sufficient and more memory efficient. Since Garf uses a much larger training dataset, we reduce the training epochs to 150, other than the 400 epochs used in Garf-mini. Both pretrainings reach over 99.5% accuracy on the validation set. Samples of segmentation results are shown in Fig. I.
Figure I: Segmentation results on a real-world object (left), Breaking Bad [42] (center) and Fantastic Breaks [18] (right).

### v2

@@ 4.2 Can Garf Generalize to Unseen Shapes?
Results and Analysis. Table 2 presents the evaluation results on the Breaking Bad dataset, where Garf achieves significant improvements over previous SOTA methods. Compared to PF++ [wang2024puzzlefusion++], Garf reduces rotation error by 82.87% and translation error by 79.83%, while achieving a CD below 0.001, indicating near-perfect reconstructions. Even more impressively, Garf-mini demonstrates exceptional generalization capability. Despite being trained only on the Everyday subset, it maintains consistent performance on the unseen Artifact subset, avoiding the performance degradation typically observed with unseen object shapes. This highlights Garf’s robust feature extraction and assembly mechanisms. Notably, Garf achieves 95.33% and 95.04% PA on the Breaking Bad dataset, approaching the theoretical maximum PA of 96.49% (Everyday subset) and 96.10% (Artifact subset) when excluding fragments smaller than 0.1% of the object volumeScanning tiny 3D fragments in real-world settings is inherently challenging; anthropologists often treat such fragments as missing parts.. On the challenging synthetic subset of Fractura, which contains unseen domain-specific objects, Garf further demonstrates superior generalization capability, outperforming all competing methods across all evaluation metrics. As shown in Fig. 4, Garf consistently produces more accurate reassemblies, particularly on the Breaking Bad Artifact subset and the Fractura synthetic fracture subset, confirming its strong generalization to unseen object shapes.
Table 2: Quantitative Results on Volume-Constrained Breaking Bad [sellan2022breaking] and Fractura Datasets. The best performance metric is highlighted in bold, while the second-best is underlined.
@@
degree $\times 10^{-2}$ % $\times 10^{-3}$
@@
Global [li2020learning] 80.50 14.60 28.70 13.00
LSTM [wu2020pq] 82.70 15.10 27.50 13.30
DGL [zhan2020generative] 80.30 13.90 31.60 11.80
Jigsaw [lu2024jigsaw] 42.19 6.85 68.89 8.22
PMTR [lee20243d] 31.57 9.95 70.60 5.56
PF++ [wang2024puzzlefusion++] 35.61 6.05 76.17 2.78
@@ 4.3 How Do Fracture Types Affect Generalization?
As scanning time significantly increases with fragment count [lamb2023fantastic], real fracture datasets are typically limited in size and diversity, making large-scale training infeasible. To address this, we investigate how fracture types impact zero-shot generalization on Fantastic Breaks and Fractura, as well as the fine-tuning performance using domain-specific synthetic fractures on Fractura.
Real Fracture on Fantastic Breaks [lamb2023fantastic]. As shown in Table 3, Garf demonstrates superior generalization to real-world fracture surfaces, achieving a remarkable 48.65% reduction in rotation error compared to PF++ [wang2024puzzlefusion++], indicating that our model effectively bridges the synthetic-to-real gap in fracture surface understanding for everyday objects.
Real Fracture on Fractura. We further evaluate performance across three fracture types. To isolate the impact of unseen objects, $\textsc{Garf}{}_{\text{LoRA}}$ is fine-tuned separately on synthetic fractures from bones, eggshells, and lithics in FracturaFine-tuning is not performed for ceramics, as its object categories closely resemble those in the Everyday subset of Breaking Bad.. Figure 5 compares reassembly performance across three fracture types in the Fractura real fracture subset. Garf outperforms competing methods on ceramics even with three missing fragments. $\textsc{Garf}{}_{\text{LoRA}}$ further improves performance by mitigating the effect of unseen objects. $\textsc{Garf}{}_{\text{LoRA}}$ generalize well to random breakage (limb bones, ceramics) and incomplete ossification (vertebrae). Unfortunately, although finetune significantly improves the performance on lithics, both Garf and $\textsc{Garf}{}_{\text{LoRA}}$ struggle on lithics, likely due to high ambiguity among flakes and the core, a well-known challenging spatial reasoning problem for anthropologists [laughlin_experimental_2010]. This unresolved challenge in Fractura offer valuable directions for future research.
Table 3: Quantitative Results on Fantastic Breaks Dataset [lamb2023fantastic]. This includes manually collected real-world objects.
@@
degree $\times 10^{-2}$ % $\times 10^{-3}$
@@ 4.5 Ablation Study
Designs of Model. We incrementally add our key design choices to the PF++ [wang2024puzzlefusion++] baseline, reporting RMSE (R/T) and PA as shown in Table 5. 1 Fracture-aware Pretraining Strategy. We replace the VAE-based pretraining used in PF++ with our fracture-aware pretraining strategy. Our strategy reduces RMSE(R) by 53.3% and RMSE(T) by 49.0%. 2 More attention on large fragments. We observe that large fragments are easier to assemble and yield more stable gradients. To leverage this, we apply surface-area-based weighted sampling and modify the attention mechanism to emphasize large fragments. This leads to substantial improvements. 3 Two-Session Flow Matching on SE(3). We design a two-session flow matching scheme: the first session performs one-step coarse pose initialization, while the second refines it via multi-step flow. This design further boosts performance during inference. More experimental results and analysis are in the supplementary materials.
Table 5: Ablation Study on Our Designs of Model.
Setups Backbone Denoiser RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
I: PuzzleFusion++ PointNet++ Diffusion 32.91 5.26 78.95
I + 1 PointNet++ Diffusion 15.36 2.68 89.40
I + 1 PointNet++ FM 15.55 2.69 89.43
I + 1 2 PointNet++ Diffusion 12.58 2.16 91.24
I + 1 2 PointNet++ FM 8.70 1.67 93.60
I + 1 2 3 PointNet++ FM 7.40 1.47 94.43
I + 1 2 3 PTv3 FM 6.68 1.34 94.77
(GARF-mini)
Figure 7: Investigation on the Number of Fragments. Garf-mini significantly outperforms PF++, with the two-session FM further boosting results for >10>10 fragments. Notably, Garf-mini also generalizes to fractures containing well over 20 fragments.
Number of Fragments. We analyze performance across varying fragment counts, as shown in Figure 7. Garf-mini consistently surpasses PF++ across all fragment counts. Our two-session FM further enhances performance on unseen objects containing more than 10 fragments. Interestingly, although Garf has only been trained on data with 2-20 fragments in Breaking Bad, it can still generalize to fractures with more than 20 pieces as shown in Figure 7.
@@ 5 Impact and Limitations
Limitations and Future Directions. Garf encounters challenges in handling fracture ambiguity, especially when fragments have subtle geometric differences or inherently ambiguous fracture surfaces. For instance, it struggles with lithic refitting in Fractura due to high ambiguity among flakes and the core, as well as fresco reconstruction in RePAIR, where erosion affects fracture surfaces [tsesmelis2025re]. Therefore, our future work will focus on: (i) Multimodal fracture reassembly, integrating geometric and texture information; (ii) Test-time policy optimization using expert feedback; (iii) Expanding the size and diversity of Fractura.
@@ 6 Related Work
Fracture Assembly. Early methods relied on explicit geometric matching with handcrafted features [luo2013co, andreadis2014facet, xu2015robust, holland2022digital], often struggling with ambiguous or incomplete geometries. The advent of the large-scale synthetic dataset Breaking Bad [sellan2022breaking] has enabled learning-based approaches to acquire robust geometric representations and assembly strategies [lu2024survey]. Jigsaw [lu2024jigsaw] jointly learns hierarchical features from global and local geometries for fracture matching and pose estimation, while SE(3)-equiv [wu2023leveraging] extracts fragment features for pose estimation. DiffAssemble [scarpellini2024diffassemble] improves performance using a diffusion model. PuzzleFusion++ [wang2024puzzlefusion++] mimics how humans solve spatial puzzles by integrating diffusion-based pose estimation with a VAE-based fragment representation and transformer-based alignment verification. However, while PuzzleFusion++ achieves SOTA results on the everyday subset, its performance degrades significantly on the Artifact subset [wang2024puzzlefusion++]. More critically, it remains unclear whether models trained on synthetic data can generalize to real-world fractures with more complex breakage patterns. To fill this gap, we identify major real-world fracture challenges and curate Fractura, a dataset capturing key complexities. To address these challenges, we propose Garf, a generalizable 3D reassembly framework for real-world fractures.
Flow Matching. Recent advances have explored flow matching (FM) across image generation [esser2024scaling], protein backbone generation [yim2023fast, huguet2024sequence, geffnerproteina], and general robot control [black2024pi0]. AssembleFlow [guo2025assembleflow] attempts to leverage FM for molecular assembly, but introduces numerical errors by approximating quaternion updates through direct addition over small time intervals during inference. While diffusion models have been widely applied to fracture assembly [scarpellini2024diffassemble, wang2024puzzlefusion++], FM provides a more natural formulation by learning geodesic flows in SE(3). We propose the first FM-based fracture assembly framework, incorporating the multi-anchor training strategy and two-session flow matching at inference time.
@@ 7 Conclusion
Acknowledgement. We gratefully acknowledge the Physical Anthropology Unit, Universidad Complutense de Madrid for access to curated human skeletons, and Dr. Scott A. Williams (NYU Anthropology Department) for the processed data samples. This work was supported in part through NSF grants 2152565, 2238968, 2322242, and 2426993, and the NYU IT High Performance Computing resources, services, and staff expertise.
@@ A.1 Fracture Simulation
(i) Bone. For elongated structures like limbs and ribs, we used Blender’s skinning and subdivision surface techniques to create realistic cylindrical hollows, replicating bone morphology. We then applied the physics-based fracture method from Breaking Bad [sellan2022breaking] to generate 2–20 fragments. The same approach was used for os coxae and vertebrae, forming the simulated subset of the bone category. (ii) Eggshell. Since scanned eggshells produce watertight solid ellipsoids, we removed 98% of the concentric volume to simulate thin shells. We then applied the same physics-based fracture method to generate realistic breakage patterns. (iii) Ceramics. Given that ceramic objects (e.g., bowls, pots, vases) closely resemble those in Breaking Bad’s everyday category, we focused on scanning real fragments and did not include a simulated subset. (iv) Lithics. As an initial feasibility test, two generalized core morphologies were repeatedly virtually knapped with some randomized variation following methods described for the dataset in [orellana2021proof] to produce core and flake combinations with varying geometries.
@@ B.1 Data Preprocessing
We preprocess the BreakingBad dataset [sellan2022breaking] to calculate the segmentation ground truth directly from meshes to reduce the computation overhead during training as described in Sec. 3.1, and there’s no need for any hyperparameters. Unlike baseline methods (Global, LSTM, and DGL) provided by the dataset and PF++ [wang2024puzzlefusion++], which samples $M=1000$ points from the mesh per fragment, we used the same setting as in Jigsaw [lu2024jigsaw] to sample $M=5000$ points per object, making all fragments have the same point density. With this sampling setting, we did not encounter any gradient explosion issues during training, as reported in FragmentDiff [xu2024fragmentdiff], which occur when sampling too many points for tiny pieces. Meanwhile, we employ the Poisson disk sampling method to ensure that the points are more uniformly distributed on the surface of the fragment. During training, standard data augmentation techniques are applied, including recentering, scaling, and random rotation.
@@ B.2 Training Recipe
We modified a smaller version of Point Transformer V3 [wu2024point] as our backbone for the segmentation pretraining, as shown in Table II, which we found to be sufficient and more memory efficient. Since Garf uses a much larger training dataset, we reduce the training epochs to 150, other than the 400 epochs used in Garf-mini. Both pretrainings reach over 99.5% accuracy on the validation set. Samples of segmentation results are shown in Fig. I.
Figure I: Segmentation results on a real-world object (left), Breaking Bad [sellan2022breaking] (center) and Fantastic Breaks [lamb2023fantastic] (right).